OpenStudy (deadshot):

Determine the zeros of f(x) = x4 - x3 + 7x2 - 9x - 18.

OpenStudy (shubhamsrg):

-1 fits in.

OpenStudy (shubhamsrg):

Got it, break 7 as 9-2 /

OpenStudy (deadshot):

so \[\pm 1\] is the answer?

OpenStudy (shubhamsrg):

nopes.

OpenStudy (shubhamsrg):

What do get after writing 7 as 9-2 ?

OpenStudy (deadshot):

\[x^4 - x^3 + 9 - 2x^2 - 9x -18\] ?

OpenStudy (shubhamsrg):

x^4 - x^3 + 9x^2 -2x^2 - 9x - 18.=0 =>(x^4 + 9x^2) - (x^3 -9x) -(2x^2 -18) =0 You see what should be the next step ?

OpenStudy (deadshot):

yeah, factor each binomial, right?

OpenStudy (shubhamsrg):

yep.

OpenStudy (deadshot):

(x^2 + 1) (x^2 + 9) for the first one?

OpenStudy (shubhamsrg):

try again.

OpenStudy (deadshot):

I'm confused

OpenStudy (amoodarya):

x1 = 2 x2 = -1 x3 = 0 + 3i x4 = 0 - 3i

OpenStudy (deadshot):

So, substitute (0 - 3i) and -1 for x^2, making it ((0 - 3i) + 9(-1))?

OpenStudy (deadshot):

so, then it would become (-3i - 9)

OpenStudy (amoodarya):

hint : divide f(x) to x^2+9 you would find your answer!

OpenStudy (deadshot):

So, divide x^4 - x^3 + 7x^2 - 9x - 18 by x^2 + 9 ?

OpenStudy (deadshot):

so it would be x^2 - x + 7 - x - 2 ?

OpenStudy (deadshot):

combine like terms, and it would be x^2 - 2x + 5 , right?

OpenStudy (anonymous):

Hmm.. Looks like you have 2 real zeros and 2 imaginary.

OpenStudy (anonymous):

Do you still need help?

OpenStudy (deadshot):

yes, I don't know what to do next

OpenStudy (deadshot):

I got it! Thanks!